# CodeAid: Evaluating a Classroom Deployment of an LLM-based Programming Assistant that Balances Student and Educator Needs

### 20 Jan 2024 v1

@@ Abstract.
Timely, personalized feedback is essential for students learning programming, especially as class sizes expand. LLM-based tools like ChatGPT offer instant support, but reveal direct answers with code, which may hinder deep conceptual engagement. We developed CodeAid, an LLM-based programming assistant delivering helpful, technically correct responses, without revealing code solutions. For example, CodeAid can answer conceptual questions, generate pseudo-code with line-by-line explanations, and annotate student’s incorrect code with fix suggestions. We deployed CodeAid in a programming class of 700 students for a 12-week semester. A thematic analysis of 8,000 usages of CodeAid was performed, further enriched by weekly surveys, and 22 student interviews. We then interviewed eight programming educators to gain further insights on CodeAid. Findings revealed students primarily used CodeAid for conceptual understanding and debugging, although a minority tried to obtain direct code. Educators appreciated CodeAid’s educational approach, and expressed concerns about occasional incorrect feedback and students defaulting to ChatGPT.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, USAbooktitle: Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, USAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Social and professional topics Computing education
@@ 1. Introduction
The recent emergence of large language models (LLMs) may offer one solution. LLM-based AI tools such as ChatGPT (OpenAI, 2022) can act as powerful coding assistants that are able to generate code from natural language descriptions. However, the rapid growth and pervasiveness of LLMs have raised concerns about their use in educational settings (Kasneci et al., 2023). This has led to some institutions banning access to tools like ChatGPT (Hsu, 2023). In computing education, in particular, concerns have been voiced regarding issues of academic integrity and student over-reliance (Becker et al., 2023). Indeed, research has shown that LLMs can generate direct solutions to almost any problem typical in introductory programming courses (Denny et al., 2023a; Finnie-Ansley et al., 2022). To develop pedagogically effective real-time support solutions for programming courses, it is necessary to implement suitable ”guardrails” that restricts the open-ended AI’s ability in generating direct solutions even when prompted. This ensures students use the AI constructively (Denny et al., 2023b).
In this paper, we present CodeAid, an LLM-based coding assistant that is designed to meet the needs of both students and educators: being helpful and technically correct, while not directly revealing code solutions. We used an iterative design approach that involved frequent requirements elicitation and feedback from the course instructor. CodeAid allowed students to (i) ask general programming questions, (ii) ask questions about the code they provide, (iii) explain the code they provide, (iv) help fix the code they provide, or (v) help write code.
@@ 3. Initial System Design and Architecture
CodeAid is an LLM-based programming assistant which aims to assist with programming assignments and reinforce concepts, similar to a teaching assistant, as outlined by Mirza et al (Mirza et al., 2019). CodeAid was designed based on prior literature, OpenAI API capabilities, instructor consultations, and pilot studies. The platform has five main features: Help Write Code and Help Fix Code for hands-on coding support, General Question and Explain Code for conceptual understanding, and Question from Code as a versatile aid covering both areas. Each feature was carefully designed to produce helpful responses while not directly generating code solutions. In this section, we focus on CodeAid’s initial design and architecture. The system went through a major update based on midterm feedback during its deployment which is described later in Section 5.
@@ 3.1.1. General Question
The most basic feature in our system was generating answers to programming questions, specifically conceptual C programming questions. When a user selects ask question, the code editor will become disabled indicating that the AI will only consider the question input. For generating the response, we used few-shot learning to generate short answers with informative explanations. See Figure 3a for more details about the prompt design for this feature. The response, displayed in Figure 2a, was limited to natural language, although sometimes included inline code (such as a function prototype), but no multi-line code.
@@ 3.1.2. Inline Code Exploration
In each of the primary five features, responses or explanations often contain C programming keywords (such as functions). To provide opportunities for learning, we displayed these keywords with a different style. With this feature, students can hover over a keyword for further exploration: generate sample code about that keyword, generate a documentation, or ask a question about the keyword. Invoking any of these three options, will generate an Inline Code Exploration response which is displayed in Figure 2e.
@@ 3.1.3. Question from Code
To simulate an experience similar to StackOverflow’s Q&A forum, we designed this feature to help students with debugging tasks or conceptual questions in a specific context. The UI for this feature looked similar to the Ask Question feature but with the added ability to provide some code as context (Figure 2b). Both the code editor and the question input became enabled when this feature was selected.
@@ 3.1.4. Help Fix Code
To help students in their code debugging tasks, students could enter their buggy code in the code editor and the intended behaviour or the problem of the code in the question input (which displayed ”Intended Behavior” as its title when this feature was selected). The initial version of the help fix code feature (Figure 2d) performed two tasks in the backend: first, it attempted to generate the correct version of the provided code based on the given description, and second, it tried to explain using bullet points what was changed and why (see Figure 3b for more details). The response interface (Figure 2d) only displayed the bullet points and not the fixed code as a way to not reveal direct code solutions.
@@ 3.1.5. Explain Code
To help students in use cases such as understanding starter code used in assignments or code taught during lectures, we designed the Explain Code feature (Figure 2c). Upon selecting this feature, the code editor would be enabled and students could paste in code they wanted to be explained. The generated output was an interface that displayed the users’ code and enabled them to hover over each line to see the detailed explanation for that line and how the line works in orchestration with the rest of the code. To do this, we used a few-shot learning approach and conditioned the model to produce a simple output structure of generating the same code but with an explanation as a specially formatted comment at the end of each line (see Figure 3c to see the structure of the prompt). This enabled us to show the explanations directly to their matching code.
@@ 3.1.6. Help Write Code
Prior work has shown LLM-based code generators to provide great starting points for programmers (Vaithilingam et al., 2022). Therefore, to assist students struggling with coding tasks, we wanted to provide ways to help them to write code without displaying any code. For that, the help write code feature (Figure 2f) required users to enter the intended behaviour of the program and generated a high-level structure of the code with sub-goals (Margulieux et al., 2012) and pseudo-code in natural language. We used few-shot learning to ensure that the generated output included information about C library functions (e.g. for memory allocation or system calls) while not including any code.
@@ 3.2. System Architecture
Figure 3. The structure of LLM prompts used in the initial version of the (a) General Question, Help Fix Code, and Explain Code features.
@@ 3.3. Prompt Design
[…]
@@ 4.1. Course Structure
The second-year course focused heavily on C programming and included topics such as shell programming, file processing, processes, signals, system calls, and basic network programming. Prerequisites of the course include a software design course taught in Java and introductory programming using Python. To accommodate the large 700-student class, the course was split into four lectures segments. Students were required to watch specific videos and complete weekly preperation exercises by a set deadline before each lecture. The course included ten lab exercises (worth 1% each), four programming assignments (A1 to A4, worth 39% in total), a midterm test (worth 10%), and a final exam (worth 40%). To incentivize responding to the weekly surveys about CodeAid, a 0.1% grade was attributed to responding to each of the ten surveys. This grade was based on completing the surveys, regardless of students’ consent to participate in our study. Course policies allowed the use of CodeAid, but explicitly forbade the use of other AI tools (such as ChatGPT) to complete any coursework. Students were also asked to cite any external sources that they used for their work. Students in the course had access to a variety of resources beyond CodeAid: recorded lecture videos, lecture notes, an online Q&A discussion board moderated by the instructors, and weekly office hours with teaching assistants and course instructors.
@@
What is being asked?
@@
How much code solution is disclosed?
@@ 4.3.4. Post-Course Anonymous Survey
Since course policy prohibited the use of code generated from AI tools like ChatGPT to complete any course work, we conducted an anonymous survey to determine whether and why students used ChatGPT. Our goal was to gain a holistic view of of students’ perceptions of AI assistants in large classes (RQ3). The survey further explored the frequency of their engagement with ChatGPT in comparison to CodeAid.
@@ 4.4. Thematic Analysis
Under each of the two dimensions, we applied an inductive approach where two researchers read through 100 randomly sampled data points together and allowed codes to emerge during the process (Bingham and Witkowsky, 2021). The process involved familiarizing themselves with the data, specifying sub-dimensions, and then creating codes for each sub-dimension. The two researchers then independently coded another 120 randomly sampled data points using the initial codebook. Next, they discussed the results from the initial coding, resolved conflicts, and further refined the codebook. During this step, they presented the codes and representative usages to the course instructor and incorporated their feedback. To enhance the generalizability and reliability of our findings, we streamlined our coding definitions. For instance, responses with minor inaccuracies were labelled as ”incorrect”. Similarly, if responses that were categorized as ”correct” failed to adequately address the query, were irrelevant, repetitive, or exceeded the scope of the course material, they were labeled as ”unhelpful”.
@@ 5.1.1. Pseudo-Code Integration
[…]
@@ 5.1.3. Stream Generation
[…]
@@ 5.1.4. Updated Prompts and OpenAI Model
[…]
@@ 5.1.5. Redesigned Help Fix Code
[…]
@@ 5.1.6. Improved Follow-Up Questions
Figure shows revised prompt design and system architecture illustrating the data flow of the updated General Question feature, highlighting the process of generating pseudo-code and relevant function documentation. The Question first gets processed by the LLM, outputting answer, code and suggested related function names. The answer is already a natural language response and is displayed to the student. Code is then processed by another LLM function to generate pseudo-code before being displayed to the student as interactive pseudo-code. Lastly, relevant function names go through a key/value, look-up table and retrieves any matching function documentation.
@@ 6.1. RQ1: Students’ Usage Patterns and Feedback
During the course, 372 students engaged with CodeAid, posing 8132 original and 1986 follow-up queries. From these, 300 students agreed that their data could be used, which left us with 7003 original queries and responses for analysis. Students, on average, inquired CodeAid 23.3 times (SD=41.2), spanning a range from 1 to a maximum of 333 questions per student. The majority of students (n=159) asked between 1 and 9 questions, followed by 80 students who asked 10 to 29 questions, 28 students who asked 30 to 49 questions, and 34 students that asked more than 50 questions. Student engagement surged approaching major assignment deadlines or exams (Figure 6b).
@@ 6.2.2. Not Displaying Direct Code Solutions
The assistant succeeded in avoiding the display of direct code solutions. In response to 43% of queries, CodeAid produced purely natural language answers which included conceptual explanations. For 24% of queries, it produced pseudo-code, of which 16% were high-level and generic example codes, while 6% were the pseudo-code that implemented a specific behavior. These specific pseudo-codes, although indirect, might have revealed the high-level ideas about implementing a particular behavior that was required in an assignment. Furthermore, when debugging code using CodeAid, it never displayed the fixed code and only recommended suggestions to fix minor syntax errors (16%) and semantic issues (8%). Similarly the Help Fix Code never generated any code. However, the initial version of General Question and Question from Code produced generic, high-level example code in 104 instances (6%) which did not directly implement any part of assignments and were similar to what students can find on websites like Stack Overflow. Finally, in 37 instances (2%) these two features generated a short code solution (1-3 lines) to a specific behavior.

### 25 Feb 2024 v2

@@ Abstract.
Timely, personalized feedback is essential for students learning programming. LLM-powered tools like ChatGPT offer instant support, but reveal direct answers with code, which may hinder deep conceptual engagement. We developed CodeAid, an LLM-powered programming assistant delivering helpful, technically correct responses, without revealing code solutions. CodeAid answers conceptual questions, generates pseudo-code with line-by-line explanations, and annotates student’s incorrect code with fix suggestions. We deployed CodeAid in a programming class of 700 students for a 12-week semester. A thematic analysis of 8,000 usages of CodeAid was performed, further enriched by weekly surveys, and 22 student interviews. We then interviewed eight programming educators to gain further insights. Our findings reveal four design considerations for future educational AI assistants: D1) exploiting AI’s unique benefits; D2) simplifying query formulation while promoting cognitive engagement; D3) avoiding direct responses while encouraging motivated learning; and D4) maintaining transparency and control for students to asses and steer AI responses.
programming education, intelligent tutoring systems, large language models, educational technology, AI assistants, AI tutoring, generative AI, class deployment, design guidelines
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642773isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Interactive systems and toolsccs: Social and professional topics Computing education
@@ 1. Introduction
The recent emergence of large language models (LLMs) may offer one solution. LLM-powered AI tools such as ChatGPT (OpenAI, 2022) can act as powerful coding assistants that generate code from natural language descriptions. However, the rapid growth and pervasiveness of LLMs have raised concerns about their use in educational settings (Kasneci et al., 2023). This has led to some institutions banning access to tools like ChatGPT (Hsu, 2023). In computing education, in particular, concerns have been voiced regarding issues of academic integrity and student over-reliance (Lau and Guo, 2023; Becker et al., 2023). Indeed, research has shown that LLMs can generate direct solutions to almost any problem typical in introductory programming courses (Denny et al., 2023a; Finnie-Ansley et al., 2022). To develop pedagogically effective real-time support solutions for programming courses, it is necessary to implement suitable ”guardrails” that restrict the open-ended AI’s ability in generating direct solutions even when prompted. This ensures students use the AI constructively (Denny et al., 2023b).
In this paper, we present CodeAid, an LLM-powered coding assistant that is designed to meet the needs of both students and educators: being helpful and technically correct, while not directly revealing code solutions. We used an iterative design approach that involved frequent requirements elicitation and feedback from the course instructor, who taught the course in which CodeAid was deployed. CodeAid allowed students to (i) ask general programming questions, (ii) ask questions about the code they provide, (iii) explain the code they provide, (iv) help fix the code they provide, or (v) help write code.
@@ 3. Initial System Design and Architecture
CodeAid is an LLM-based programming assistant which aims to assist with programming assignments and reinforce concepts, similar to a teaching assistant, as outlined by Mirza et al (Mirza et al., 2019). CodeAid was designed based on prior literature, OpenAI API capabilities, instructor consultations, and pilot studies. The platform has five main features: Help Write Code and Help Fix Code for hands-on coding support, General Question and Explain Code for conceptual understanding, and Question from Code as a versatile assistance covering both areas. Each feature was carefully designed to produce helpful responses while not directly generating code solutions. In this section, we focus on CodeAid’s initial design and architecture. The system went through a major update based on midterm feedback during its deployment which is described later in Section 5.
@@ 3.1.1. General Question
[Uncaptioned image] The most basic feature in our system was generating answers to programming questions, specifically conceptual C programming questions. When a user selects General Question, the code editor will become disabled indicating that the AI will only consider the question input. For generating the response, we used few-shot learning to generate short answers with informative explanations. See Figure 3a for more details about the prompt design for this feature. The response, displayed in Figure 2a, was limited to natural language, although sometimes included inline code (such as a function prototype), but no multi-line code.
@@ 3.1.2. Inline Code Exploration
[Uncaptioned image] In each of the primary five features, responses or explanations often contain C programming keywords (such as functions). To provide opportunities for learning, we displayed these keywords in a different style. With this feature, students can hover over a keyword for further exploration: generate sample code about that keyword, generate documentation, or ask a question about the keyword. Invoking any of these three options, will generate an Inline Code Exploration response which is displayed in Figure 2e.
@@ 3.1.3. Question from Code
[Uncaptioned image] To simulate an experience similar to StackOverflow’s Q&A forum, we designed the Question from Code feature to help students with debugging tasks or conceptual questions in a specific context. The UI for this feature looked similar to the General Question feature but with the added ability to provide some code as context (Figure 2b). Both the code editor and the question input became enabled when this feature was selected.
@@ 3.1.4. Help Fix Code
[Uncaptioned image] To help students in their code debugging tasks, students could enter their buggy code in the code editor and the intended behaviour or the problem of the code in the question input (which displayed ”Intended Behavior” as its title when this feature was selected). The initial version of the Help Fix Code feature (Figure 2d) performed two tasks in the backend: first, it attempted to generate the correct version of the provided code based on the given description, and second, it tried to explain using bullet points what was changed and why (see Figure 3b for more details). The response interface (Figure 2d) only displayed the bullet points and not the fixed code as a way to not reveal direct code solutions.
@@ 3.1.5. Explain Code
[Uncaptioned image] To help students in use cases such as understanding starter code used in assignments or code taught during lectures, we designed the Explain Code feature (Figure 2c). Upon selecting this feature, the code editor would be enabled and students could paste in code they wanted to be explained. The generated output was an interface that displayed the users’ code and enabled them to hover over each line to see the detailed explanation for that line and how the line works in orchestration with the rest of the code. To do this, we used a few-shot learning approach and conditioned the model to produce a simple output structure of generating the same code but with an explanation as a specially formatted comment at the end of each line (see Figure 3c to see the structure of the prompt). This enabled us to show the explanations directly to their matching code.
@@ 3.1.6. Help Write Code
[Uncaptioned image] Prior work has shown LLM-based code generators to provide great starting points for programmers (Vaithilingam et al., 2022). Therefore, to assist students struggling with coding tasks, we wanted to provide ways to help them to write code without displaying any code. For that, the Help Write Code feature (Figure 2f) required users to enter the intended behaviour of the program and generated a high-level structure of the code with sub-goals (Margulieux et al., 2012) and pseudo-code in natural language. We used few-shot learning to ensure that the generated output included information about C library functions (e.g. for memory allocation or system calls) while not including any code.
@@ 3.2. System Architecture
Figure 3. The structure of LLM prompts used in the initial version of General Question, Help Fix Code, and Explain Code.
@@ 3.3. Prompt Design
Our prompt engineering consisted of mainly few-shot training, in which at least one input/output example was provided for each prompt. We carefully designed prompts to ensure responses (i) follow structured, and easy-to-parse templates, (ii) are technically correct, and (iii) use a style, tone, and level of technicality that is both helpful and not overwhelming for students. For example, in refining the Help Fix Code feature, we moved from generating a simple bullet-pointed suggestions to first generating the correct code, followed by suggested fixes. This significantly improved suggestion quality and accuracy. Similarly, in the updated system, we adopted a pseudo-code style that balances between not overly revealing the code’s syntax and not being too close to natural language, which might be too long and overwhelming. Furthermore, the initial system used the [Uncaptioned image] GPT-3 model, but was updated to the [Uncaptioned image] GPT-3.5 model in the updated system. The structure of the prompts used in the General Question, Help Fix Code, and Explain Code features of the initial version of CodeAid is illustrated in the Figure 3.
@@ 4.1. Course Structure
The second-year course focused heavily on C programming, and included topics such as shell programming, file processing, processes, signals, system calls, and basic network programming. Prerequisites of the course include a software design course taught in Java and introductory programming using Python. To accommodate the large 700-student class, the course was split into four lectures segments. Students were required to watch specific videos and complete weekly preparation exercises by a set deadline before each lecture. The course included ten lab exercises (worth 1% each), four programming assignments (A1 to A4, worth 39% in total), a midterm test (worth 10%), and a final exam (worth 40%). To incentivize responding to the weekly surveys about CodeAid, a 0.1% grade was attributed to responding to each of the ten surveys. This grade was based on completing the surveys, regardless of students’ consent to participate in our study. Course policies allowed the use of CodeAid, but explicitly forbade the use of other AI tools (such as ChatGPT) to complete any coursework. Students were also asked to cite any external sources that they used for their work. Students in the course had access to a variety of resources beyond CodeAid: recorded lecture videos, lecture notes, an online Q&A discussion board moderated by the instructors, and weekly office hours with teaching assistants and course instructors.
@@
What are students asking from CodeAid?
@@
How much is CodeAid directly revealing the solution?
@@ 4.3.4. Post-Course Anonymous Survey
Since course policy prohibited the use of code generated from AI tools like ChatGPT to complete any course work, we conducted an anonymous survey to determine whether and why students used ChatGPT. Our goal was to gain a holistic view of students’ perceptions of AI assistants in large classes (RQ3). The survey further explored the frequency of their engagement with ChatGPT in comparison to CodeAid.
@@ 4.4. Thematic Analysis
Under each of the two dimensions, we applied an inductive approach where two researchers read through 100 randomly sampled data points together and allowed codes to emerge during the process (Bingham and Witkowsky, 2021). The process involved familiarizing themselves with the data, specifying sub-dimensions, and then creating codes for each sub-dimension. The two researchers then independently coded another 120 randomly sampled data points using the initial codebook. Next, they discussed the results from the initial coding, resolved conflicts, and further refined the codebook. During this step, they presented the codes and representative usages to the course instructor and incorporated their feedback. To enhance the generalizability and reliability of our findings, we streamlined our coding definitions. For instance, responses with minor inaccuracies were labelled as ”incorrect”. Similarly, if responses that were categorized as ”correct” failed to adequately address the query, were irrelevant, repetitive, or exceeded the scope of the course material, they were labelled as ”unhelpful”.
@@ 5.1.1. Pseudo-Code Integration
As a way to provide more comprehensive responses and to increase engagement, we decided to add pseudo-code generation to most of the features (Figure 4b). We used pseudo-code as a form of scaffolding, offering a simplified and structured outline of a program, which serves as a bridge to actual coding without directly revealing the code itself (Lee and Ko, 2015). To enable this functionality, we updated our LLM prompts for all features to ask the model to generate code. However, instead of showing this code to the user, we sent it to a new LLM function that generated the pseudo-code. For line-by-line explanations, the new LLM function also included an explanation following each line in the form: [Uncaptioned image]. This was then parsed and only the pseudo-code was rendered, while the explanation would be displayed whenever the user hovers over that particular line.
@@ 5.1.3. Stream Generation
To address concerns over response delays, we incorporated OpenAI’s stream generation mechanism for CodeAid’s responses to provide immediate feedback. To achieve this, we developed a specific markup for each of the features to enable parsing the partially generated response as it was being streamed. This enabled our system to immediately start displaying responses after users clicked the generate button. For single-line components such as generating a short summary for the provided code, or responding with a single-line answer to an asked question, our markup used the following format: [Uncaptioned image]. For multi-line components such as code parts, we used two tokens, [Uncaptioned image] and [Uncaptioned image]. Furthermore, in situations where the output of one LLM function was sent to another LLM function in series (e.g. to generate the final pseudo-code from code), our API did not show the generated code, but displayed (and updated while streaming) the number of lines of code that were generated.
@@ 5.1.4. Updated Prompts and OpenAI Model
We improved the General Question and Question from Code features to provide more detailed and thorough answers by refining the few-shot prompting examples. To enhance the overall quality, accuracy, and reliability of CodeAid’s responses, we upgraded the AI model from [Uncaptioned image] to the more advanced [Uncaptioned image] model. We also updated the prompts to not respond to questions that were not relevant to C programming.
@@ 5.1.5. Redesigned Help Fix Code
We redesigned the feature based on feedback so that it would highlight the specific lines that require modifications, deletions, or additions. To achieve this, we used a data flow which is displayed in Figure 5. (1) Pre-processing: The buggy code is stripped of any comments and is then reformatted with a standard style. (2) Generating Fixed Code: An LLM function tries to generate the fixed version of the buggy code based on the provided intended behaviour or error message. This step also generates a paragraph of changes similar to the initial version of the feature which is immediately streamed to the client and displayed. (3) Matching Lines: A simple static code analyzer matches each line of the fixed code to the original buggy code. (4) Annotating Buggy Code: The buggy code is then annotated with three labels: [Uncaptioned image], [Uncaptioned image], or a new empty line with [Uncaptioned image]. (5) Explaining Annotations: The annotated buggy code and the fixed code are sent to another LLM function that adds explanations to each of the changed, removed, and added labels. Eventually, the explained changes and annotated code is streamed to the client for rendering and displaying the highlighted lines and on-hover interactions as displayed in Figure 4a.
@@ 5.1.6. Improved Follow-Up Questions
Figure shows revised prompt design and system architecture illustrating the data flow of the updated General Question feature, highlighting the process of generating pseudo-code and relevant function documentation. The Question first gets processed by the LLM, outputting answer, code and suggested related function names. The answer is already a natural language response and is displayed to the student. Code is then processed by another LLM function to generate pseudo-code before being displayed to the student as interactive pseudo-code. Lastly, relevant function names go through a key/value, look-up table and retrieve any matching function documentation.
@@ 6.1. RQ1: Students’ Usage Patterns and Feedback
During the course, 372 students engaged with CodeAid, posing 8132 original and 1986 follow-up queries. From these, 300 students agreed that their data could be used, which left us with 7003 original queries and responses for analysis. Students, on average, inquired CodeAid 23.3 times (SD=41.2), spanning a range from 1 to a maximum of 333 questions per student. The majority of students (n=159) asked between 1 and 9 questions, followed by 80 students who asked 10 to 29 questions, 28 students who asked 30 to 49 questions, and 34 students who asked more than 50 questions. Student engagement surged approaching major assignment deadlines or exams (Figure 6b). The dataset of queries from consented students is available as supplementary material with this paper, offering a valuable resource for those interested in advancing AI educational tools.
@@ 6.2.2. Not Displaying Direct Code Solutions
The assistant succeeded in avoiding the display of direct code solutions. In response to 43% of queries, CodeAid produced purely natural language answers which included conceptual explanations. For 24% of queries, it produced pseudo-code, of which 16% were high-level and generic example codes, while 6% were the pseudo-code that implemented a specific behavior. These specific pseudo-codes, although indirect, might have revealed the high-level ideas about implementing a particular behavior that was required in an assignment. Furthermore, when debugging code using CodeAid, it never displayed the fixed code and only recommended suggestions to fix minor syntax errors (16%) and semantic issues (8%). Similarly, the Help Fix Code never generated any code. However, the initial version of General Question and Question from Code produced generic, high-level example code in 104 instances (6%) which did not directly implement any part of assignments and were similar to what students can find on websites like Stack Overflow. Finally, in 37 instances (2%) these two features generated a short code solution (1-3 lines) to a specific behavior.